Shift Summary, 10/09/99 Day Shift

Tony Barker (tonyb@cuhep.Colorado.EDU)
Sat, 9 Oct 1999 14:44:46 -0600 (MDT)

Shift Summary for Day Shift, 10/09/99

On Shift: Hogan Nguyen, Tony Barker

Hours of beam available at all intensities: 0:00
Hours of beam available at nominal intensity: 0:00
Hours of E832 data (special runs not included): 0:00
Hours of special runs (laser scans, muon runs...): 0:00

Accounting of lost time: TeV He purification continues.

========================================================================

Accesses during shift: Hall was open at start; search and secure
completed 11:30.

Controlled access at 14:30 to reset Etot
LV power strip.

========================================================================

Overview of shift: Not much going on.

Elliott, Jason H, and Sydney notice an apparently
inefficient strip in Etot, which then turns out to
be due to a tripped LV power supply causing Etot
to have failed for half the CsI. However, most
of the channels float high, so the they appear to
have HCC bits on even though they are not working.
Problem turns out to be a tripped power strip;
cause unknown.

========================================================================
Items to Follow up on Next Shift

Keep an eye on the Etot LV power alarm to make sure power
strip doesn't trip off again.